OpenStudy (anonymous):

Find the center-radius form of the equation of the circle. center (-2, 0), radius = square root of 2

OpenStudy (zzr0ck3r):

(x-x_1)^2 +(y-y_1)^2 = r^2 x_1 = -2, y_1 = 0, and r = sqrt(2)

OpenStudy (zzr0ck3r):

can you do it now?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh yes i got it now thank you!!!

OpenStudy (zzr0ck3r):

what you get for an answer

OpenStudy (anonymous):

(x + 2)^2 + y^2=sqrt2 is that right

Directrix (directrix):

@caroesfea (x + 2)^2 + y^2=(sqrt2)^2 = 2 (x + 2)^2 + y^2 = 2 --> Don't forget to square the radius

OpenStudy (anonymous):

@Directrix thank you!

Directrix (directrix):

Glad to help.